# BLIPs: Bayesian Learned Interatomic Potentials

### v1

@@ 4 Experiments
We demonstrate the effectiveness of BLIP across a range of simulation-based chemistry tasks. First, as an illustrative example, we use BLIP to model a controlled dynamical system of charged particles interacting via Coulomb forces in a bounded 3D space. We then apply BLIP to two challenging MLIP tasks: (i) training from scratch using the equivariant PaiNN architecture (Schütt et al.,, 2021) on a data-scarce, out-of-distribution ammonia (NH3) system, and (ii) fine-tuning the pretrained ORB v3 model (Rhodes et al.,, 2025) on silica glass (SiO2) with 699 atoms. In both cases, BLIP improves uncertainty calibration and predictive performance.
@@ 4.1 Charged Particle System
As an illustrative example, we begin by training BLIP to learn the dynamics of five charged particles confined in a three-dimensional box, where their interactions evolve according to Coulomb’s law. This setup serves as a simplified example to test the model’s ability to learn and perform uncertainty quantification in a simpler scenario than simulation-based chemistry.
Dataset and Metrics:
We model a system composed of 5 particles, each carrying a positive or negative charge and associated with a position and velocity in 3D space. The dataset is provided by (Fuchs et al.,, 2020), and we use the same split for training and testing. The task is to predict the particle positions after $1000$ timesteps, given their initial positions and velocities. We train the model using the standard mean squared error (MSE) loss, augmented with a KL divergence term for BLIP. We evaluate performance using three metrics: mean square error (MSE), negative log-likelihood (NLL), and the continuous ranked probability score (CRPS). Further dataset and metrics details are provided in Section B.
Implementation Details:
We benchmark two architectures for this task: a standard Graph Neural Network (GNN) (Gilmer et al.,, 2017) and the Equivariant Graph Neural Network (EGNN) proposed by (Satorras et al.,, 2021) using the same network hyperparameters and training protocol as in (Satorras et al.,, 2021). For the GNN baseline, edge attributes are also encoded via a linear layer before being used in message passing, as we noticed a performance improvement. For each neural architecture, we compare BLIP with different variants. First, we train the deterministic version of the models to establish a baseline for predictive accuracy. We then evaluate two Bayesian versions of the same architectures for UQ: Monte Carlo (MC) Dropout (Gal and Ghahramani,, 2016), using different dropout probabilities; and Deep Ensembles (Lakshminarayanan et al.,, 2017) with four ensemble members. For BLIP, we employ an invariant inference network, using an MLP consisting of two linear layers of size $64$ with swish activations, used for both the message and update coefficients. The MLP work directly on the invariant features without message-passing. The message inference network head takes as input the edge attribute $a_{ij}$ together with $\mathbf{h}^{0}_{i}$ and $\mathbf{h}^{0}_{j}$, whereas the update inference network head only takes $\mathbf{h}^{0}$. Additional architecture and training details are provided in Section C.
Results:
Table 1: Future position mean square error (MSE) in the N-body system. Averages are computed over 4 random initialisation seeds.
@@ 4.2 Learning Interatomic Potentials in Data-Scarse and Out-Of-Distribution Regimes
Machine Learning Interatomic Potentials (MLIPs) are surrogate models that approximate the potential energy surface (PES) of atomic systems. Once the PES is learned, atomic forces can be directly computed as the negative gradient of the energy with respect to atomic positions. Accurate MLIP models enable highly efficient atomic system simulations, paving the way for various applications, such as catalyst design, adsorbents for recycling CO2 from the atmosphere, and efficient solar cells, among others. This experiment focuses on the study of the Ammonia (NH3 ) molecule for assessing the model’s robustness to out-of-distribution (OOD) data. This is a common challenge in materials science, where training data may be biased toward equilibrium or near-equilibrium configurations.
Dataset and metrics:
We use the Ammonia dataset introduced by (Tan et al.,, 2023), which consists of a small number of NH3 molecular geometries sampled across a range of energy configurations. The training set includes 78 geometries corresponding to near-equilibrium configurations with energies in the range of 0 to 5 kcal/mol. The test set, in contrast, contains 129 out-of-distribution (OOD) geometries with energies spanning up to 70 kcal/mol, thereby evaluating model generalisation far from the training distribution. This setup enables us to assess both predictive accuracy and the reliability of uncertainty estimates in a realistic low-data, high-OOD regime. We evaluate performance using three metrics: mean absolute error (MAE), expected calibration error (ECE), and the Spearman correlation. Further dataset and metrics details are provided in Section B.
@@
We adopt the (equivariant) PaiNN architecture (Schütt et al.,, 2021) as our machine learning interatomic potential. To ensure a fair comparison, we use the same hyperparameter settings, training protocol, and evaluation procedure as described in (Tan et al.,, 2023). We train the model by explicitly enforcing energy conservation, computing atomic forces as the negative gradient of the predicted potential energy with respect to atomic positions, and weighting the energy/forces terms accordingly to Tan et al., (2023). For BLIP, we employ a compact invariant posterior network consisting of two linear layers of size $64$ with swish activations, used for both the message and update coefficients. The message inference network head takes as input the square norm of the positions $\|\bm{r}_{i}-\bm{r}_{j}\|^{2}$, whereas the update inference network head only takes an embedding of the element atomic number $Z$. Additional architecture and training details are provided in Section C.
@@ Results:
Table 2: Energy and Forces mean absolute error (MAE) in the Ammonia system. Averages are computed over 4 random initialisation seeds. Baseline models’ results are from Tan et al., (2023).
@@
Table 3: Expected calibration error (ECE) for Forces [kcal/mol/Å] in the Ammonia system for OOD data. Averages are computed over 4 random initialisation seeds. Baseline models’ results are from Tan et al., (2023).
Model ECE ($\downarrow$)
Deep Ensemble $0.108$
Evidential Regression $0.120$
Mean Variance Estimate $0.135$
Gaussian Mixture Model $0.112$
BLIP (ours) $\mathbf{0.091}^{\pm 0.006}$
As shown in Table 2, BLIP achieves the best MAE performance across both energy and forces targets, outperforming all baseline models from (Tan et al.,, 2023). For energy prediction, BLIP achieves an MAE of $11.51\pm 0.75$, slightly better than the Deep Ensemble ($11.71$) and substantially better than other probabilistic methods. On force prediction, which is a more challenging task, BLIP shows a notable improvement with an MAE of $41.93\pm 3.38$, significantly lower than both the Deep Ensemble and other baselines. These results demonstrate that training an MLIP with BLIP yields superior predictive accuracy on both energy and force targets, without modifying the base architecture or requiring multiple models, as is the case with Deep Ensembles.
@@
Regarding uncertainty estimation, Table 4.2 reports the expected calibration error (ECE) for the different models. ECE measures how well the estimated probabilities match the observed probabilities, which intuitively means that if a model predicts to be confident $n\%$ of the time, it should be correct approximately $n\%$ of the time. From the table, it is clear that BLIP is the most calibrated model, outperforming even the Deep Ensemble method. Additionally, we analyse the linear relationship between predicted uncertainty (variance) and squared error (Figure 2), which is particularly relevant for active learning. Our model demonstrates a strong linear correlation (Spearman correlation of $0.863\pm 0.024$), indicating its reliability in identifying structures where the MLIP predictions are likely inaccurate.
4.3 Finetuning Pretrained Interatomic Potentials
@@
The final set of experiments focuses on fine-tuning a pretrained MLIP. In the current era of increasing data and compute scale, training universal models from scratch is often impractical without large computional resources. Instead, it is common to fine-tune existing pretrained models. Among recent advancements, we select the ORB v3 Direct 20OMat model (Rhodes et al.,, 2025), a state-of-the-art direct force field that is not rotationally-equivariant. We fine-tune this model on a dataset composed of various silica (SiO2) structures extracted from long and diverse molecular dynamics simulations. Each configuration contains 699 atoms in total, including 233 silicon and 466 oxygen atoms. This dataset presents a realistic and challenging benchmark, as the formation of large bulk structures makes DFT calculations computationally expensive.
Dataset and metrics:
The silica glass (SiO2) dataset introduced by (Tan et al.,, 2023) consists of system snapshots extracted from diverse, long molecular dynamics trajectories. These trajectories capture a range of physical conditions, including temperature equilibration, cooling, uniaxial tension, compression, and more (see Tan et al., (2023) for details on dataset generation). We fine-tune models using varying numbers of training samples ($32$, $128$, and $1024$) and evaluate performance on a held-out test set of $373$ samples. Importantly, no structurally similar configurations appear in both the training and test sets, ensuring a robust evaluation. An example configuration from the dataset is shown in Figure 3(a). This experimental setup allows us to assess both predictive accuracy under varying data regimes and the reliability of uncertainty estimates in a realistic application. We report performance using three metrics: mean absolute error (MAE), expected calibration error (ECE), and Spearman correlation. Further details on the dataset and evaluation metrics can be found in Section B.
@@
We fine-tune the non-conservative ORB v3 model, originally trained on a subset of ab-initio molecular dynamics data from OMAT 24 (Barroso-Luque et al.,, 2024). We select ORB v3 due to its strong performance among state-of-the-art MLIPs and its low inference latency, benefiting from its non-conservative nature, which eliminates the need for automatic differentiation during force computation. All models are fine-tuned for 50 epochs using a batch size of 8, in float32-high precision, with the AdamW optimiser and cosine scheduler. The loss function minimises the Huber loss (as in Rhodes et al., (2025)) for both energy and force predictions. For BLIP, we employ a compact posterior network composed of four linear layers with hidden size 64 and swish activations for both the message and update functions. The edge network takes as input a concatenation of the pairwise distance vector (normalised to unit length) and 20 Gaussian radial basis functions applied to the raw edge distance. The node network processes atomic number embeddings. Finally, dropout probabilities $p$ are predicted and transformed into multiplicative noise coefficients using the expression p/(1p)p/(1-p). Additional architectural and training details are provided in Section C.
@@ Results:
[…]
@@
Figure 3(b) shows the results of fine-tuning the ORB v3 model on the Silica dataset. For comparison, we include in the caption the best-performing model reported in the literature on this dataset, a Deep Ensemble Equivariant PaiNN (4 members) from (Tan et al.,, 2023). Notably, this model yields a substantially higher error than even the pretrained ORB v3 model (pretrained), highlighting the potential advantages of universal MLIPs. We then fine-tuned ORB v3 both with (w/) and without (w/o) BLIP. For inference with the BLIP fine-tuned models, we evaluated predicted forces using two approaches: (i) the sample mean computed from 100 stochastic forward passes (sample mean, Algorithm 3), and (ii) a single forward pass using the unperturbed weights (Gal and Ghahramani,, 2016; Blundell et al.,, 2015), i.e., the maximum a posteriori estimate (MAP estimate, Algorithm 4). We also fine-tuned a Deep Ensemble of 4 members for direct comparison.
@@
The results demonstrate that BLIP is both accurate and efficient. Using the sample mean or the more efficient MAP estimate, BLIP outperforms all ORB v3 baseline models and the pretrained PaiNN ensemble. Remarkably, BLIP with MAP inference achieves better accuracy than deepensembling the fine-tuned models, at a fraction of the computational cost, since it requires storing and fine-tuning only a single model. For uncertainty quantification, we evaluate ECE and Spearman correlation on the $1024$-sample setup. BLIP achieves an ECE of $0.032\pm 0.002$, while Deep Ensemble performs significantly worse with $0.247$. When examining the correlation between predicted variance and squared error, crucial for active learning, BLIP slightly underperform, with a Spearman correlation of $0.602\pm 0.002$ versus $0.676$ for Deep Ensemble.
Overall, the results highlight that BLIP provides accurate calibration and reliable uncertainty estimates for guiding model refinement. More importantly, fine-tuning with BLIP achieves higher predictive accuracy than both deterministic and ensemble baselines, while maintaining the same inference cost as a standard deterministic model.
@@ 5 Conclusions
We introduced BLIP, a scalable and architecture-agnostic variational Bayesian framework for principled uncertainty quantification in (equivariant) message passing neural networks. Based on Adaptive Variational Dropout (AVD), BLIP can be used for both training from scratch and fine-tuning pretrained MPNNs. Unlike deep ensembles, it maintains the same inference-time efficiency as standard deterministic models for predictive tasks, while enabling well-calibrated uncertainty estimates through stochastic sampling. We demonstrated the effectiveness of BLIP across a range of scientific tasks, including particle prediction and machine-learned interatomic potentials. BLIP achieves remarkable performance in prediction and uncertainty quantification for MLIPs when trained in data-scarce and OOD regimes. Notably, we showed that fine-tuning pretrained MLIPs with BLIP consistently enhances model performance, outperforming both deterministic and ensemble-based baselines. In addition, the uncertainty estimates produced by BLIP are well-calibrated and strongly correlated with model error, making them particularly useful for downstream applications such as active learning and error detection. These results highlight the potential of BLIP to serve as a drop-in tool for improving the performance of MLIPs, or more generally MPNNs, and provide them with useful uncertainty estimates at no additional computational cost.
Impact Statement
This paper presents work aimed at advancing the field of machine learning for atomistic modeling, with the potential to accelerate scientific discovery in chemistry and the biological sciences. As machine learning–based atomistic modeling continues to evolve, we recognize that there are numerous potential societal consequences of this work. While we do not feel that any specific issues need to be highlighted, we emphasize the importance of ongoing discussions on ethical considerations and best practices to ensure these technologies are developed and applied in ways that benefit society as a whole.
Acknowledgement
D. Coscia acknowledges the support provided by PRIN "FaReX - Full and Reduced order modelling of coupled systems: focus on non-matching methods and automatic learning" project. We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy). D. Coscia conducted the project when visiting the University of Amsterdam.
Supplementary Materials
[…]
[sections] Table of Contents
[…]
[sections]l1

### v2

@@ 4 Experiments
We demonstrate the effectiveness of BLIP across a range of simulation-based chemistry tasks. First, as an illustrative example, we use BLIP to model a system of charged particles interacting via Coulomb forces. We then apply BLIP to three challenging MLIP tasks: (i) training from scratch using the equivariant PaiNN architecture (Schütt et al., 2021) on a data-scarce, out-of-distribution ammonia (NH3) system, and fine-tuning the pretrained ORB v3 model (Rhodes et al., 2025) on (ii) the silica glass (SiO2) with 699 atoms, and (iii) the MATPES (Kaplan et al., 2025) PBE dataset using active learning. In both cases, BLIP improves uncertainty calibration and predictive performance.
@@ 4.1 Charged Particle System
Table 1: Results for the N-body system. Averages are computed over 4 random initialisation seeds.
@@
We use the Coloumbs’ N-body particle benchmark (Fuchs et al., 2020) as an initial example to test BLIP’s uncertainty quantification abilities. The system consists of 5 charged (positive or negative) particles, each equipped with a position and velocity in 3D space. The task is to predict the particle positions after $1000$ timesteps, given their initial positions and velocities. We use the same split for training and testing as Satorras et al. (2021), and trained the models using the standard mean squared error loss, augmented with a KL divergence term for BLIP. Further dataset and metrics details are provided in Section B.
Implementation Details:
We benchmark two architectures for this task: a standard Graph Neural Network (GNN) (Gilmer et al., 2017) and the Equivariant Graph Neural Network (EGNN) proposed by (Satorras et al., 2021) using the same network hyperparameters and training protocol as in (Satorras et al., 2021). We then evaluate two Bayesian versions of the same architectures for UQ: Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016), using different dropout probabilities $p$; and Deep Ensembles (Lakshminarayanan et al., 2017) with four ensemble members. For BLIP, a compact invariant posterior network for both the message and update function coefficients is used. Additional architecture and training details are provided in Section C.
Results:
@@ 4.2 Learning Interatomic Potentials in Data-Scarse and Out-Of-Distribution Regimes
Table 2: Expected calibration error for Forces in the Ammonia system for OOD data. Averages are computed over 4 random initialisation seeds. Baseline models’ results are from Tan et al. (2023).
Model ECE ($\downarrow$) $\rho$ ($\uparrow$)
Deep Ensemble $0.108$ $0.825$
Evidential Regression $0.120$ $0.779$
Mean Variance Estimate $0.135$ $0.693$
Gaussian Mixture Model $0.112$ $\mathbf{0.904}$
BLIP (ours) $\mathbf{0.091}^{\pm 0.006}$ $0.863^{\pm 0.024}$
This experiment focuses on the study of the Ammonia (NH3) molecule for assessing the model’s robustness in a realistic low-data, high out-of-distribution (OOD) regime. This is a common challenge in materials science, where training data may be biased toward equilibrium or near-equilibrium configurations. We use the Ammonia dataset introduced by (Tan et al., 2023), which consists of a small number of NH3 molecular geometries sampled across a range of energy configurations. The training set includes 78 geometries corresponding to near-equilibrium configurations with energies in the range of 0 to 5 kcal/mol. The test set, in contrast, contains 129 OOD geometries with energies spanning up to 70 kcal/mol, thereby evaluating model generalisation far from the training distribution. We evaluate performance using three metrics: mean absolute error (MAE), expected calibration error (ECE), and the Spearman correlation. Further dataset and metrics details are provided in Section B.
@@
We adopt the (equivariant) PaiNN architecture (Schütt et al., 2021) as our machine learning interatomic potential. To ensure a fair comparison, we use the same hyperparameter settings, training protocol, and evaluation procedure as described in (Tan et al., 2023). We train the model by explicitly enforcing energy conservation, computing atomic forces as the negative gradient of the predicted potential energy with respect to atomic positions, and weighting the energy/forces terms accordingly to Tan et al. (2023). For BLIP, a compact invariant posterior network for both the message and update function coefficients is used. Additional architecture and training details are provided in Section C.
@@ Results:
Table 3: Energy and Forces mean absolute error (MAE) in the Ammonia system. Averages are computed over 4 random initialisation seeds. Baseline models’ results are from Tan et al. (2023).
@@
As shown in Table 3, BLIP achieves the best MAE performance across both energy and forces targets, outperforming all baseline models from (Tan et al., 2023). For energy prediction, BLIP achieves an MAE of $11.51\pm 0.75$, slightly better than the Deep Ensemble ($11.71$) and substantially better than other probabilistic methods. On force prediction, which is a more challenging task, BLIP shows a notable improvement with an MAE of $41.93\pm 3.38$, significantly lower than both the Deep Ensemble and other baselines. These results demonstrate that training an MLIP with BLIP yields superior predictive accuracy on both energy and force targets, without modifying the base architecture or requiring multiple models, as is the case with Deep Ensembles. Regarding uncertainty estimation, Table 4.2 reports the expected calibration error (ECE) for the different models. ECE measures how well the estimated probabilities match the observed probabilities, which intuitively means that if a model predicts to be confident $n\%$ of the time, it should be correct approximately $n\%$ of the time. From the table, it is clear that BLIP is the most calibrated model, outperforming even the Deep Ensemble method. Additionally, we analyse the linear relationship between predicted uncertainty (variance) and squared error (Figure 2), which is particularly relevant for active learning. Our model demonstrates a strong linear correlation (Spearman correlation of $0.863\pm 0.024$), indicating its reliability in identifying structures where the MLIP predictions are likely inaccurate.
4.3 Finetuning Pretrained Interatomic Potentials on SiO2
@@
This experiment focuses on fine-tuning a pretrained MLIP on SiO2. In the current era of increasing data and compute scale, training universal models from scratch is often impractical without large computional resources. Instead, it is common to fine-tune existing pretrained models. Among recent advancements, we select the ORB v3 Direct 20OMat model (Rhodes et al., 2025), a state-of-the-art direct force field that is not rotationally equivariant. We fine-tune the silica glass (SiO2) dataset introduced by (Tan et al., 2023), which consists of system snapshots extracted from diverse, long molecular dynamics trajectories. Each configuration contains 699 atoms in total, including 233 silicon and 466 oxygen atoms. This dataset presents a realistic and challenging benchmark, as the formation of large bulk structures makes DFT calculations computationally expensive. An example configuration from the dataset is shown in Figure 3(a). We fine-tune models using varying numbers of training samples ($32$, $128$, and $1024$) and evaluate performance on a held-out test set of $373$ samples. We report performance using three metrics: mean absolute error (MAE), expected calibration error (ECE), and Spearman correlation. Further details on the dataset and evaluation metrics can be found in Section B.
@@
We fine-tune the non-conservative ORB v3 model, originally trained on a subset of ab-initio molecular dynamics data from OMAT 24 (Barroso-Luque et al., 2024). We select ORB v3 due to its strong performance among state-of-the-art MLIPs and its low inference latency, benefiting from its non-conservative nature, which eliminates the need for automatic differentiation during force computation. We fine-tuned ORB v3 both with (w/) and without (w/o) BLIP. We also fine-tuned a Deep Ensemble of 4 members for direct UQ comparison. Additional architectural and training details are provided in Section C.
@@ Results:
Figure 3: (a) Visualisation of the silica glass structure. (b) Forces mean absolute error (MAE) in the Silica system across different training sizes for finetuned ORB v3 direct 20OMAT. Averages are computed over 4 random initialisation seeds, and the standard error of the mean is reported as an error bar. The best baseline model is a Deep Ensemble Equivariant PaiNN from Tan et al. (2023) trained on the $1024$ configuration (same split), achieving MAE $6.58\text{ kcal/mol/Å}$.
@@
Figure 3(b) shows the results of fine-tuning the ORB v3 model on the Silica dataset. For comparison, we include in the caption the best-performing model reported in the literature on this dataset, a Deep Ensemble Equivariant PaiNN (4 members) from (Tan et al., 2023). Notably, this model yields a substantially higher error than even the pretrained ORB v3 model (pretrained), highlighting the potential advantages of universal MLIPs. For inference with the BLIP fine-tuned models, we evaluated predicted forces using two approaches: (i) the sample mean computed from 100 stochastic forward passes (sample mean, Algorithm 3), and (ii) a single forward pass using the unperturbed weights (Gal and Ghahramani, 2016; Blundell et al., 2015), i.e., the maximum a posteriori estimate (MAP estimate, Algorithm 4).
@@
The results demonstrate that BLIP is both accurate and efficient. Using the sample mean or the more efficient MAP estimate, BLIP outperforms all ORB v3 baseline models and the pretrained PaiNN ensemble. Remarkably, BLIP with MAP inference achieves better accuracy than deepensembling the fine-tuned models, at a fraction of the computational cost, since it requires storing and fine-tuning only a single model. For uncertainty quantification, we evaluate ECE and Spearman correlation on the $1024$-sample setup. BLIP achieves an ECE of $0.032\pm 0.002$, while Deep Ensemble performs significantly worse with $0.247$. When examining the correlation between predicted variance and squared error, crucial for active learning, BLIP slightly underperform, with a Spearman correlation of $0.602\pm 0.002$ versus $0.676$ for Deep Ensemble. Overall, the results highlight that BLIP provides accurate calibration and reliable uncertainty estimates for guiding model refinement. More importantly, fine-tuning with BLIP achieves higher predictive accuracy than both deterministic and ensemble baselines, while maintaining the same inference cost as a standard deterministic model.
4.4 Active Learning Fine-Tuning of Pretrained Interatomic Potentials
In this experiment, we fine-tune the ORB v3 Direct 20OMat model on the MATPES PBE dataset (Kaplan et al., 2025), which comprises roughly 400,000 carefully sampled structures from 281 million molecular dynamics snapshots, covering 16 billion atomic environments.
Table 4: Energy and Forces mean absolute error (MAE) in the MATPES PBE dataset. Averages and standard deviations are computed over 3 random initialization seeds.
Model MAE ($\downarrow$)
Energy [meV/atom] Forces [meV/Å]
w/o BLIP $11.83^{\pm 1.24} 11$$10.84^{\pm 0.57}$
w/ BLIP + random $96.65^{\pm 0.09}$ $105.12^{\pm 0.06}$
w/ BLIP + uncertainties $\mathbf{91.13}^{\pm 1.14}$ $\mathbf{103.85}^{\pm 0.14}$
To efficiently improve model accuracy, we employ an Active Learning (AL) strategy. In this framework, BLIP uncertainties are used to identify structures where the pretrained model is most uncertain. These high-uncertainty structures are then computed at the reference DFT level and used to iteratively fine-tune the model. By repeating this cycle, the model progressively learns from the most informative data points, maximizing performance gains while minimizing computational cost. We assess model performance using the mean absolute error (MAE) for both forces and energies, and we track the evolution of the validation loss throughout the AL iterations. Detailed descriptions of the dataset and evaluation metrics are provided in Section B, while implementation details in Section C.
Figure 4: Validation loss for different models during each active learning step, calculated as the sum of mean absolute errors (MAE) for energy and forces. Results are averaged over 3 independent training runs. Energy is reported in eV and forces in eV/Å.
Results:
The validation loss in Figure 4 exhibits a clear and consistent order of convergence across models. Fine-tuning with BLIP using uncertainty-guided selection of structures (w/ BLIP + uncertainties), where structures are chosen based on the variance in predicted energies from a pool, achieves the fastest and most stable convergence to lower error. Fine-tuning with BLIP on randomly selected structures (w/ BLIP + random) converges to a worse optimum, but still outperforms the model without BLIP (w/o BLIP), which shows the poorest performance. Importantly, additional active learning steps further amplify the performance gap between uncertainty-guided and random selection strategies. This trend is also reflected in the test errors for both forces and energies, which follow the same ordering (Table 4). These results demonstrate that fine-tuning with BLIP consistently improves model quality, and that leveraging BLIP uncertainties for active learning not only accelerates convergence but also systematically enhances predictive accuracy.
@@ 5 Conclusions
We introduced BLIP, a scalable and architecture-agnostic variational Bayesian framework for principled uncertainty quantification in (equivariant) message passing neural networks. BLIP can be used for both training from scratch and fine-tuning pretrained MPNNs. Unlike deep ensembles, it maintains the same inference-time efficiency as standard deterministic models for predictive tasks, while enabling well-calibrated uncertainty estimates through stochastic sampling. We demonstrated the effectiveness of BLIP across a range of scientific tasks, including particle prediction and machine-learned interatomic potentials. BLIP achieves remarkable performance in prediction and uncertainty quantification for MLIPs when trained in data-scarce and OOD regimes. Notably, we showed that fine-tuning pretrained MLIPs with BLIP consistently enhances model performance, outperforming both deterministic and ensemble-based baselines. In addition, the uncertainty estimates produced by BLIP are highly calibrated and strongly correlated with model error, making them particularly useful for downstream applications such as active learning and error detection. These results highlight the potential of BLIP to serve as a drop-in tool for improving the performance of MLIPs, more generally MPNNs, and provide them with useful uncertainty estimates.
Appendix